OpenStudy (anonymous):

a student in greece discovers a pottery bowl that contains 62% of its original amount of c-14. find the age of the pottery bowl to the nearest year

OpenStudy (wolf1728):

First of all we need the half life of carbon 14 which is 5,730 years Next we need the formula for the time that has passed: http://www.1728.org/halflife.htm time=half-life * [log(bgng amt/endg amt) / log(2)] time=5,730 * [log(100/62)/ log(2)] time = 5,730 * 0.2076083105 / 0.3010299957 time = 3,951.75 years (This agrees with the calculator on that page.)

OpenStudy (mrnood):

YOU NEED TO BE AWARE THAT C14 DATING IS NOT ACCURATE TO 1 YEAR This is a very misleading question

OpenStudy (wolf1728):

It is also misleading because you wouldn't use carbon dating on inorganic substances such as pottery.

OpenStudy (anonymous):

ancient pottery may contain straw

OpenStudy (wolf1728):

Also, wikipedia says the halflife of Carbon 14 is 5,730 years +- 40 years. I was just going by the mathematics of it.

OpenStudy (anonymous):

thank you! ive been stuck on that question for decades lol

OpenStudy (wolf1728):

Yes, those half life questions can be VERY difficult. I was glad to help out. :-)